OpenStudy (anonymous):

what is the simplified form of 3(2x-5)>21

OpenStudy (asnaseer):

ok, first divide both sides by 3 - what do you end up with?

OpenStudy (anonymous):

3 cant go into 2, nor 5. but can into 21 "7" times

OpenStudy (asnaseer):

if you have an equation like:\[a(b+c)=d\]then if you divide both sides by a, what you get is:\[\frac{1}{\cancel{a}}\times \cancel{a}(b+c)=\frac{1}{a}\times d\]\[b+c=\frac{d}{a}\]

OpenStudy (asnaseer):

so, in your case, after dividing both sides by 3 we get:\[2x-5\gt7\]

OpenStudy (asnaseer):

can you solve the rest of this now?

OpenStudy (anonymous):

ugh no. i give up!

OpenStudy (anonymous):

thanks anyway for tryin to help.

OpenStudy (asnaseer):

why do you say "ugh no" to this?

OpenStudy (anonymous):

because i don't get it.

OpenStudy (asnaseer):

you don't get how we divided both sides by 3?

OpenStudy (anonymous):

no ecause i don see how u can do 3 into those numbers.without makin it intoa decimal

OpenStudy (asnaseer):

ok, lets take a simple example to illustrate this. lets say we have: 3 * 7 = 21 now we can divide both sides by 3 to get: 7 = 21/3 = 7 agreed?

OpenStudy (asnaseer):

does that make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, now lets replace the 7 with (5+2), and we get: 3(5+2) = 21 this is basically the same equation - agreed?

OpenStudy (anonymous):

yess

OpenStudy (asnaseer):

so you need to think of everything in the braces as ONE SINGLE ITEM. 3(5+2) = 21 divide both sides by 3 to get: 5+2 = 21/3 = 7

OpenStudy (asnaseer):

does that help you understand this better?

OpenStudy (anonymous):

yes.. it does THANKS:)!!!

OpenStudy (asnaseer):

gr8! - so do you think you can solve from where we left off now?

OpenStudy (asnaseer):

last step above was:\[2x-5\gt7\]

OpenStudy (asnaseer):

so, for example, the next step would be to add 5 to both sides